OpenStudy (anonymous):

A waffle recipe calls for different amounts of mix, depending on the number of servings. For 4 servings, Jordan needs 3 cups of mix and for 12 servings; she needs 9 cups of mix. This shows a direct variation between the number of servings (x) and the amount of mix (y). Using the formula y =kx, solve the formula for k and then give the value of k.

OpenStudy (anonymous):

just substitute x servings and y mix in. 3 = k 4 9 = k 12 what is k?

OpenStudy (anonymous):

is it 108

OpenStudy (anonymous):

noooooo o_O k = 3/4 k = 9/12

OpenStudy (anonymous):

oooo

OpenStudy (anonymous):

is either 1 of them 2 my answers

OpenStudy (anonymous):

both are the same. 0.75 :P